If a 23.7 mL solution of H2SO4 has a concentration of 0.100 M, what is its pH? Also, how
many grams of H2SO4 are in the solution?
1
Expert's answer
2020-05-03T15:25:08-0400

Solution.

"V(H_2SO_4)=23.7mL=0.0237L;"

"M(H_2SO_4)=0.100M;"

"[H_3O^+]=[H_2SO4]=0.100M;"

"pH=-lg[H_3O^+]=-lg(0.100)=1;"

"\nu(H_2SO_4)=\dfrac{m}{Molecular weight};\implies m=\nu(H_2SO_4)\sdot Molecular weight;"

"Molecular weight=98g\/mol;"

"\nu(H_2SO_4)=V(H_2SO_4)\sdot M; \nu=0.0237L\sdot0.100mol/L=0.00237mol;"

"m=0.00237mol\sdot 98g/mol=0.232g;"

Answer: "pH=1;"

"m=0.232g."